OpenStudy (anonymous):

Write the point-slope form of the equation of the line passing through the point ( -2,4 ) with a slope of -3

OpenStudy (aakashsudhakar):

x1 = -2; y1 = 4; m = -3 Point-Slope Equation form: y - y1 = m(x - x1) Therefore, the equation is... y - 4 = -3(x + 2)

OpenStudy (anonymous):

Thank you